OpenStudy (anonymous):

Find all real numbers (x) (if any) such that F(g(x)) = g(f(x)). Where F(x) = x^2+1 AND g(X) = 1 - X

OpenStudy (ranga):

Find f(g(x)) and g(f(x)) first. Equate them and then solve for x.

OpenStudy (anonymous):

Okay I put \[x ^{2} + 1 = 1 - x\] . Then, Added the single X. and Subtracted the single (+) 1. to end up at \[x ^{3} = 0\]. So... For the function of F = All real numbers. and for G = All real numbers above (greater than) 1?

OpenStudy (anonymous):

When you're adding an X-> You do X^{2} right? not 2x. .... Ummm The part where I added the single X. Did I do that right or did i do that wrong?

OpenStudy (ranga):

Need clarification on something. Does the problem have uppercase and lowercase F and f? Because then it will be two different functions.

OpenStudy (anonymous):

Lower case

OpenStudy (ranga):

f(x) = x^2 + 1 g(x) = 1 - x f(g(x)) = (1-x)^2 + 1 = 1 + x^2 - 2x + 1 = x^2 - 2x + 2 g(f(x)) = 1 - (x^2 + 1) = 1 - x^2 - 1 = -x^2 f(g(x)) = g(f(x)) x^2 - 2x + 2 = -x^2 2x^2 - 2x + 2 = 0 x^2 - x + 1 = 0 solve for x.

OpenStudy (ranga):

f(x) = x^2 + 1 To find f(g(x)), replace x with g(x): f(g(x)) = (g(x))^2 + 1 = (1-x)^2 + 1 (as done above)

OpenStudy (anonymous):

So, do i have to factor at \[x ^{2}-x+1 = 0\]

OpenStudy (ranga):

Use the quadratic formula.

OpenStudy (ranga):

If you find the discriminant, b^2 - 4ac, you will notice it is negative. Therefore, there are no real values of x that will satisfy the condition.

OpenStudy (anonymous):

So since it cannot be factored, there is no real number that can satisfy X. So the answer is No real numbers?

OpenStudy (ranga):

The answer is "No REAL Numbers" but not because it cannot be factored. When we use the quadratic formula we get a negative number inside the square root. That means the equation has complex roots and no real roots.

OpenStudy (anonymous):

Thank You for your help :)

OpenStudy (ranga):

You are welcome.